# Vision-Language-Action Models: Concepts, Progress, Applications and Challenges

### v1

@@ 5.2 Future Roadmap
The future of VLA models lies at the intersection of increasingly powerful multimodal foundations, agentic reasoning, and embodied continual learning. Over the next decade, we anticipate several converging trends that will propel VLAs from capable but narrow task specialists toward the core of truly generalist robotic intelligence.
  1. 1.
    Multimodal Foundation Models as the “Cortex.” Today’s VLAs typically couple a vision-language backbone with task-specific policy heads. Tomorrow, we expect a single, massive multimodal foundation model—trained on web-scale image, video, text, and affordance data—to serve as a shared perceptual and conceptual “cortex.” This foundation will encode not only static scenes but also dynamics, physics, and common-sense world knowledge, enabling downstream action learners to tap into a unified representational substrate rather than reinventing basic perceptual skills for every robot or domain.
  2. 2.
    Agentic, Self-Supervised Lifelong Learning. Rather than static pretraining, future VLAs will engage in continual, self-supervised interaction with their environments. Agentic frameworks—where the model generates its own exploration objectives, hypothesizes outcomes, and self-corrects via simulated or real rollouts—will drive rapid skill acquisition. By formulating internal sub-goals (“learn to open drawers,” “map furniture affordances”) and integrating reinforcement-style feedback, a VLA-driven humanoid could autonomously expand its capabilities over years of deployment, much like a human apprentice.
  3. 3.
    Hierarchical, Neuro-Symbolic Planning. To scale from low-level motor primitives to high-level reasoning, VLAs will adopt hierarchical control architectures. A top-level language-grounded planner (perhaps an LLM variant fine-tuned for affordance reasoning) will decompose complex instructions (“prepare a cup of tea”) into sequences of sub-tasks (“fetch kettle,” “fill water,” “heat water,” “steep tea bag”). Mid-level modules will translate these into parameterized motion plans, and low-level diffusion or transformer-based controllers will generate smooth, compliant trajectories in real time. This neuro-symbolic blend ensures both the interpretability of structured plans and the flexibility of learned policies.
  4. 4.
    Real-Time Adaptation via World Models. Robustness in unstructured settings demands that VLAs maintain an internal, predictive world model—an up-to-date simulation of objects, contacts, and agent dynamics. As the robot acts, it will continuously reconcile its predictions with sensor feedback, using model-based corrective actions when discrepancies arise (e.g., slipping grasp). Advances in differentiable physics and video-to-state encoders will make these world models both accurate and efficient enough for on-board, real-time use. Cross-Embodiment and Transfer Learning: The era of training separate VLAs for each robot morphology will give way to embodiment-agnostic policies. By encoding actions in an abstract, kinematic-agnostic space (e.g., “apply grasp force at these affordance points”), future VLAs will transfer skills seamlessly between wheeled platforms, quadrupeds, and humanoids. Combined with meta-learning, a new robot can bootstrap prior skills with only a few minutes of calibration data. Safety, Ethics, and Human-Centered Alignment As VLAs gain autonomy, built-in safety and value alignment become non-negotiable. Future systems will integrate real-time risk estimators—assessing potential harm to humans or property before executing high-risk maneuvers—and seek natural language consent for ambiguous situations. Regulatory constraints and socially aware policies will be baked into the VLA stack, ensuring that robots defer to human preferences and legal norms.
As illustrated in Figure 18, the future of VLA-based robotics lies in the integration of three foundational components: Vision-Language Models (VLMs), VLA architectures, and agentic AI systems. Consider “Eva,” a generalist humanoid assistant operating in a household. At the perception layer, Eva’s foundation VLM interprets multimodal inputs by segmenting visual scenes into discrete object-level representations, predicting affordances (e.g., graspable, fragile), and simulating dynamic behaviors through an internal world model. This VLM layer enables high-level visual understanding grounded in language semantics and physical properties. Upon receiving a user command such as “Eva, clean the coffee spill and water the plants,” the VLA module activates. This core architecture combines tokenized language inputs and sensory feedback to perform hierarchical task planning. A high-level planner decomposes the instruction into actionable subtasks (e.g., locate cloth, wipe spill, retrieve watering can), which are then converted into motion trajectories via a mid-level policy module. These plans are passed to a low-level diffusion-policy controller, responsible for generating smooth, physics-aware joint movements tailored to the robot’s embodiment. Complementing these is Eva’s agentic AI module, which supports continual learning and adaptation. When confronted with unexpected challenges—like a sticky stain—Eva invokes an internal self-improvement loop, running real-time simulated variations to refine its wiping strategy without human supervision. Safety and alignment are ensured through human-aware policies: proximity sensors, real-time monitoring, and verbal confirmations before high-risk actions. Overnight, Eva performs autonomous review of performance logs, refining sub-policies via simulated rollouts. Together, this VLM-VLA-agentic triad marks a significant leap toward embodied AGI. It enables robots like Eva to perceive, plan, act, adapt, and safely coexist with humans, ultimately transforming how intelligent systems interact with real-world environments in robust, interpretable, and human-aligned ways.
@@ 6 Conclusion
In this comprehensive review, we systematically evaluated the recent developments, methodologies, and applications of VLA models published over the last three years. Our analysis began with the foundational concepts of VLAs, defining their role as multimodal systems that unify visual perception, natural language understanding, and action generation in physical or simulated environments. We traced their evolution and timeline, detailing key milestones that marked the transition from isolated perception-action modules to fully unified, instruction-following robotic agents. We highlighted how multimodal integration has matured—from loosely coupled pipelines to transformer-based architectures that enable seamless coordination between modalities.
@@
Next, we examined tokenization and representation techniques, focusing on how VLAs encode visual and linguistic information, including action primitives and spatial semantics. We explored learning paradigms, detailing the datasets and training strategies—from supervised learning and imitation learning to reinforcement learning and multimodal pretraining—that have shaped VLA performance. In our section on adaptive control and real-time execution, we addressed how modern VLAs are optimized for dynamic environments, discussing policies that support latency-sensitive tasks. We then categorized major architectural innovations, surveying over 50 recent VLA models. This included advancements in model design, memory systems, and interaction fidelity. We further studied training and efficiency strategies, including parameter-efficient methods like LoRA, quantization, and model pruning, alongside acceleration techniques such as parallel decoding and hardware-aware inference. Our analysis continued with real-world applications of VLA models, showcasing their deployment across six domains: humanoid robotics, autonomous vehicles, industrial automation, healthcare, agriculture, and augmented reality (AR) navigation. Each application was reviewed with examples of model performance, domain-specific challenges, and generalizability.
@@
In addressing challenges and limitations, we focused on five core areas: real-time inference, multimodal action representation and safety, bias and generalization, system integration and compute constraints, and ethical deployment. We proposed potential solutions drawn from current literature, including model compression, cross-modal grounding, domain adaptation, and agentic learning frameworks. Finally, our discussion and future roadmap articulated how the convergence of VLMs, VLA architectures, and agentic AI systems is steering robotics toward artificial general intelligence (AGI). This review provides a unified understanding of VLA advancements, identifies unresolved challenges, and outlines a structured path forward for developing intelligent, embodied, and human-aligned agents.
Acknowledgement
@@
This work was supported by the National Science Foundation and the United States Department of Agriculture, National Institute of Food and Agriculture through the “Artificial Intelligence (AI) Institute for Agriculture” Program under Award AWD003473, and AWD004595, Accession Number 1029004, ”Robotic Blossom Thinning with Soft Manipulators”.
@@
ChatGPT and Perplexity were utilized to enhance grammatical accuracy and refine sentence structure; all AI-generated revisions were thoroughly reviewed and edited for relevance. Additionally, ChatGPT-4o was employed to generate realistic visualizations.
Appendix Table
The following appendix tables present a comprehensive overview of recent developments and challenges in VLA models. Table 3 systematically compares state-of-the-art VLA methodologies, their application domains, and key innovations across robotics, autonomous systems, and embodied AI platforms. This comparative summary highlights core architectural advances, deployment contexts, and technical contributions—providing valuable insight into the evolving landscape of generalist and task-specific VLA models. Meanwhile, Table 4 presents a structured synthesis of the major technical and practical challenges facing VLA model deployment, alongside potential solutions and their expected impact. This includes limitations such as real-time inference constraints, multimodal integration issues, and ethical concerns, with proposed resolutions ranging from architectural innovations to scalable training techniques and cross-modal alignment strategies. Together, these tables serve as a detailed reference for researchers, developers, and practitioners aiming to understand both the current capabilities and outstanding barriers in VLA-based intelligent systems.

### v2

@@ 5.2 Future Roadmap
The future of VLA-based systems are expected to evolved at the intersection of increasingly capable multi-modal foundations, agentic reasoning, and embodied continual learning. Over the next decade, we anticipate several converging trends that will propel VLAs from capable yet brittle task specialists toward dependable, generalist robotic intelligence. However, this trajectory will be shaped by persistent constraints/limitations highlighted earlier: (i) real-time inference bottlenecks in closed-loop control, (ii) incomplete multi-modal action representations and weak safety assurance, (iii) dataset bias and grounding failures under distribution shift, (iv) integration complexity across perception-memory-reasoning-control, (v) high compute and energy demands that hinder edge deployment, and (vi) robustness, transparency, and ethical concerns in open-world settings. To address these issues holistically, Figure 19 summarizes a system-level research roadmap, while Figure 18 provides an intuitive conceptual illustration of how VLMs, VLA architectures, and agentic AI modules may co-evolve toward embodied AGI in robotics.
Future RoadmapEfficientDeploymentReliable &Safe IntelligenceUnified Systems& GovernanceParameter-efficientVLA backbonesAnytime / early-exitinferenceCompact actiontokenization & chunkingOn-device caching& episodic memoryHardware-awarecompilation (edge)Robust multimodalgrounding (uncertainty)Calibrated abstention& safe fallbackWorld models:physical & causal predictionConstraint-aware control(shields/MPC)Verification & runtimesafety monitorsUnified 2D–temporal–3DrepresentationsCross-embodimenttransfer & adaptationSim-to-real curricula& domain randomizationEvaluation beyond success(safety, energy, recovery)Governance: privacy,bias audits, accountability
Figure 19: Future research roadmap for Vision-Language-Action (VLA) models (Fig. 19). Dashed boxes denote prioritized directions addressing the core limitations of VLA models identified in this work: (i) Efficient Deployment (latency, compute, energy), (ii) Reliable & Safe Intelligence (grounding, uncertainty, safety assurance), and (iii) Unified Systems & Governance (2D-temporal-3D integration, transfer, evaluation, and responsible deployment).
Multi-modal foundation models as the “cortex” for embodied perception:
Today’s VLA stacks often rely on a vision-language backbone coupled to task-specific policy heads, which limits reuse of general knowledge and increases retraining cost across domains. A plausible next step is a unified multi-modal foundation model trained on web-scale image, video, text, and interaction/affordance traces to function as a shared “cortex” that encodes not only static semantics but also dynamics, contact priors, and commonsense physical knowledge [295, 289, 272]. Such a cortex can reduce failure modes caused by shallow correlations by grounding language in object-centric representations and persistent scene structure [206]. As emphasized in Figure 18, this foundation-model cortex would enable robots to segment environments into actionable entities (objects, regions, affordances) and provide stable semantic anchors to downstream planners and controllers. Yet, to prevent overconfidence and hallucinated grounding, these models must incorporate calibrated uncertainty and evidence-linked reasoning to ensure that perception-driven plans remain verifiable under occlusion, clutter, and ambiguous instructions [144].
Agentic, self-supervised, lifelong learning and continual adaptation:
A defining limitation of current VLAs is their static nature: policies trained once are deployed unchanged, despite operating in non-stationary environments. Future VLAs should adopt agentic learning loops where models propose exploration objectives, hypothesize outcomes, and self-correct through simulated and real rollouts, enabling continual skill growth over months or years [43, 90]. This direction is expected to mitigate distribution shift, dataset bias, and long-horizon brittleness by allowing models to adapt over time; however, continual policy updates introduce new risks. In particular, repeated online or incremental learning can overwrite previously acquired competencies (catastrophic forgetting), induce unintended behavioral regressions, and increase susceptibility to noisy, adversarial, or unintended environmental feedback that may corrupt the learned policy [292, 263]. Therefore, lifelong learning must be coupled with replay and safety-aware updates, modular adapters, and verification-informed policy revisions [121]. Within the roadmap of Figure 19, this agentic lifelong learning paradigm falls naturally at the intersection of Reliable & Safe Intelligence and Unified Systems & Governance, where continual learning is treated as a controlled, auditable lifecycle process rather than an ad hoc fine-tuning step.
Hierarchical, neuro-symbolic planning for scalability and interpretability:
Scaling from low-level motor primitives to long-horizon objectives requires explicit hierarchy [261, 265]. Next-generation VLA systems will likely use language-grounded planners (LLM-style modules fine-tuned for affordances and constraints) that decompose goals into structured sub-tasks, followed by mid-level skill policies and low-level controllers that ensure compliant motion [94, 253]. This neuro-symbolic blend helps bridge integration complexity by imposing interfaces that are easier to debug, monitor, and certify [201]. Importantly, hierarchy also enables selective verification: high-level plans can be checked for constraint violations (unsafe steps, forbidden regions) [192, 71], while low-level trajectories can be shielded by control barrier functions, MPC, and runtime safety monitors [188, 196]. These components align with Figure 19 under the pillar of Reliable & Safe Intelligence, where safety is enforced through both planning-time constraints and execution-time guards rather than post hoc evaluation alone [101, 274].
Real-time adaptation via world models and physical/causal reasoning:
Robust deployment in unstructured settings demands that VLAs maintain internal predictive models of objects, contacts, and dynamics. World models that forecast near-term state transitions and failure likelihoods can support counterfactual evaluation (“if I push here, what collides?”) and rapid corrective actions when reality deviates from expectation (e.g., slipping grasp, unexpected friction) [203]. This capability is central to safe manipulation, navigation, and human-robot interaction, where small errors compound quickly [156]. Yet, world models must be efficient enough for on-board use and consistent with multi-sensor evidence. Thus, a major future direction is hardware-aware, memory-efficient predictive modeling such as temporal token compression [267, 216], event-driven state updates [250, 226], and hybrid physics-learning models that combine differentiable physical simulators with learned dynamics to enable control-relevant update rates [102, 52]. In Figure 19, these needs appear jointly in the Efficient Deployment pillar (real-time constraints) and the Reliable & Safe Intelligence pillar (physics/causality for grounded decision-making).
Efficiency and scalability: bridging generality with edge deployment:
A central barrier to VLA adoption remains the mismatch between the computational footprint of large multi-modal backbones and the latency/energy constraints of closed-loop control [268, 238]. Future VLAs should prioritize parameter-efficient designs (structured sparsity, low-rank adaptation, modular experts) that preserve generalization while reducing inference cost [65, 197]. Beyond training-time efficiency, anytime/early-exit policies can allocate computation adaptively, ensuring that safety-critical steps retain high fidelity while routine steps use cheaper pathways [41, 274]. Equally important is action-space efficiency: compact action tokenization and chunked control representations shorten autoregressive horizons, enabling higher control rates without sacrificing temporal smoothness [171]. These model-level choices must be paired with hardware-aware compilation across GPU/NPU/edge accelerators, including quantization-aware scheduling and memory-optimized attention kernels [65, 168]. Compute-aware caching and episodic memory further reduce redundant forward passes and improve responsiveness in long-horizon tasks [27, 133]. Collectively, these directions operationalize the Efficient Deployment pillar of Figure 19 and directly address the compute/energy limitations emphasized earlier.
Cross-embodiment transfer and morphology-agnostic skill representations:
The approach of training separate VLAs for each robot morphology is unlikely to scale. A key future theme is embodiment-agnostic policy learning, where skills are expressed in abstract action spaces (e.g., contact goals, affordance-point manipulation, task-space constraints) that transfer across wheeled platforms, quadrupeds, and humanoids [275, 118]. Meta-learning and few-shot calibration can allow rapid bootstrapping on new robots with minutes of data rather than weeks of training [35]. This direction also mitigates dataset bias by enforcing invariances across embodiments and environments, but it demands standardized representations, common interfaces, and reproducible evaluation protocols [118, 275]. In Figure 19, this embodiment-agnostic skill learning paradigm sits under Unified Systems & Governance, linking architectural unification with principled transfer and benchmarking.
Evaluation beyond task success: safety, recovery, and resource-aware metrics:
Progress in VLA requires measurement that reflects deployment realities. Task success alone does not clearly represent failure severity, temporal inconsistency, unsafe near-misses, and energy inefficiency. Future benchmarks should quantify safety violations, uncertainty calibration, recovery behavior, temporal coherence, energy consumption, and downstream utility under human constraints [274, 78]. Moreover, evaluations should report compute budgets, dataset composition, and deployment conditions to enable fair comparisons and diagnose bias-driven gains [215, 273, 118]. Such measurement is not merely scientific hygiene: it is the foundation for auditing and governance, and it determines whether a system can be responsibly deployed at scale [274, 275]. This motivation underlies the evaluation branch of Figure 19 and complements the conceptual deployment narrative in Figure 18.
Safety, ethics, and human-centered alignment as first-class design objectives:
As VLAs gain autonomy, built-in safety and value alignment become critical. Future systems should integrate real-time risk estimators that assess potential harm before executing high-risk actions, request natural-language confirmation under ambiguity, and maintain transparent logs for accountability [274, 118]. Privacy-aware sensing, bias audits, and human-in-the-loop oversight must be embedded into the lifecycle, especially for assistive robotics and safety-critical autonomy [284, 236]. Regulatory-aligned evaluation protocols and standardization efforts will be essential to translate VLA advances into trusted real-world systems [169, 294, 235]. This governance perspective is explicitly captured in Figure 19 and is implicitly reflected by the socially aware humanoid setting illustrated in Figure 18.
Cross-cutting themes: continual learning, failure recovery, interaction, and control fidelity:
Across all pillars in Figure 19, several cross-cutting themes are expected to shape the next decade of VLA research. First, continual and lifelong learning must be safe, auditable, and resistant to forgetting, enabling long-term adaptation without destabilizing deployed systems [290]. Second, failure detection and recovery should be treated as first-class capabilities, incorporating introspective monitoring, uncertainty-aware perception, and structured recovery behaviors when execution deviates from expected outcomes [116, 259]. Third, improving the precision and reliability of action generation remains critical: while VLA-based planners enable flexible, language-conditioned decision making, their trajectory accuracy and control stability currently lag behind conventional analytical approaches such as model predictive control, sampling-based motion planning, and feedback-linearized controllers. As a result, hybrid architectures that combine VLA-driven high-level planning with classical or learned low-level controllers are expected to play a central role in achieving both semantic flexibility and control-level precision. Finally, human alignment and interaction require mechanisms for intent clarification, shared autonomy, and explainable action rationales, supporting trust and usability across diverse real-world settings [111, 110].
In summary, Figure 19 emphasizes that closing the gap between laboratory demonstrations and robust real-world deployment will require coordinated advances in efficiency, safety, data and generalization, system integration, evaluation, and governance. Complementarily, Figure 18 illustrates how these advances may converge toward generalist embodied agents across multiple platforms including mobile robots, manipulators, assistive systems, and humanoids where multi-modal perception, hierarchical planning, continual adaptation, and human-aligned safety are integrated within a unified intelligence stack. Addressing these directions collectively is expected to transform VLAs from promising research prototypes into dependable, broadly applicable embodied systems rather than solutions limited to any single robot morphology.
@@ 6 Conclusion
In this comprehensive review, we systematically evaluated the recent developments, methodologies, and applications of Vision-Language-Action (VLA) models published over the last three years. Our analysis began with the foundational concepts of VLAs, defining their role as multi-modal systems that unify visual perception, natural language understanding, and action generation in physical or simulated environments. We traced their evolution and timeline, detailing key milestones that marked the transition from isolated perception-action modules to fully unified, instruction-following robotic agents. We highlighted how multi-modal integration has matured from loosely coupled pipelines to transformer-based architectures that enable seamless coordination between modalities.
@@
Next, we examined tokenization and representation techniques, focusing on how VLAs encode visual and linguistic information, including action primitives and spatial semantics. We explored learning paradigms, detailing the datasets and training strategies from supervised learning and imitation learning to reinforcement learning and multi-modal pretraining that have shaped VLA performance. In ’adaptive control and real-time execution’ section, we discussed how modern VLAs are optimized for dynamic environments, analyzing policies that support latency-sensitive tasks. We then categorized major architectural innovations, surveying over 50 recent VLA models. This discussion included advancements in model design, memory systems, and interaction fidelity. We further studied strategies for training efficiency improvement, including parameter-efficient methods like LoRA, quantization, and model pruning, alongside acceleration techniques such as parallel decoding and hardware-aware inference. Our analysis of real-world applications highlighted both the promise and current limitations of VLA models across six domains: humanoid robotics, autonomous vehicles, industrial automation, healthcare, agriculture, and augmented reality (AR) navigation. Across these settings, VLAs demonstrated strong capabilities in high-level semantic reasoning, instruction-following, and task generalization, particularly in structured or partially controlled environments. However, their effectiveness was often constrained by real-time inference latency, limited robustness under environmental variability, and reduced precision in long-horizon or safety-critical control when compared to conventional analytical planning and control pipelines. Moreover, application-specific adaptations and extensive data curation were frequently required to achieve reliable performance, underscoring challenges in scalability and deployment. These findings suggest that while VLAs are well-suited for semantic decision making and flexible task specification, hybrid architectures that integrate VLA reasoning with classical or learned low-level controllers remain essential for practical, real-world operation.
@@
In addressing challenges and limitations, we focused on five core areas: real-time inference, multi-modal action representation and safety, bias and generalization, system integration and compute constraints, and ethical deployment. We proposed potential solutions drawn from current literature, including model compression, cross-modal grounding, domain adaptation, and agentic learning frameworks. Finally, our discussion and future roadmap articulated how the convergence of VLMs, VLA architectures, and agentic AI systems is steering robotics toward artificial general intelligence (AGI). This review provides a unified understanding of VLA advancements, identifies unresolved challenges, and outlines a structured path forward for developing intelligent, embodied, and human-aligned agents in the future.
Funding Declaration
@@
This work was supported in part by the National Science Foundation (NSF) and the United States Department of Agriculture (USDA), National Institute of Food and Agriculture (NIFA), through the “Artificial Intelligence (AI) Institute for Agriculture” program under Award Numbers AWD003473 and AWD004595, and USDA-NIFA Accession Number 1029004 for the project titled “Robotic Blossom Thinning with Soft Manipulators.” Additional support was provided through USDA/NIFA Grant Number 2024-67022-41788, Accession Number 1031712, under the project “ExPanding UCF AI Research To Novel Agricultural EngineeRing Applications (PARTNER).
@@
ChatGPT and Perplexity were utilized to enhance grammatical accuracy and refine sentence structure; all AI-generated revisions were thoroughly reviewed and edited for relevance.